Form a polynomial f(x) with real coefficients having the given degree and zeros.
Degree: 3; zeros: -2 and 3 + i.

Answer by ewatrrr(24785)

(x+2)[x-(3+i)][x - (3-i)] = (x+2)[(x-3)^2 - i^2] = (x+2)(x^2 - 6x + 9 + 1)
(x+2)(x^2 - 6x + 10)
Multiply to finish